OpenStudy (anonymous):

All- Does the graph of this quadratic equation open upward or downward? Explain your answer. f(x) = -(-x²-8x-3)

OpenStudy (anonymous):

 ax^2+bx+c when a>0 it goes up and here it's -(-x^2-8x-3)=x^2+8x+3 so a=1 which is higher than 0

OpenStudy (radar):

f(x)=x^2 +8x + 3 Note that the coefficient of the x^2 term is a positive 1 As Tomas.A pointed out if the coefficient (a) is > 0 (positive) it opens upward.